OpenStudy (anonymous):

what is the value of -9+(2^3-3^2)

OpenStudy (bakonloverk):

-7 2^3=8 3^2=6 8-6=2 -9+2= -7

OpenStudy (anonymous):

-9+(8-9)=-10

OpenStudy (bakonloverk):

oh did math wrong yea it is -10